Mr. Hernandez had to interview a total of 63
households in his assignment. He has already
finished 28. What percentage of the households
in his assignments has he finished? (Round
your answer to the nearest tenth of a percent.)

Answer by stanbon(75887)

Mr. Hernandez had to interview a total of 63
households in his assignment. He has already
finished 28. What percentage of the households
in his assignments has he finished? (Round
your answer to the nearest tenth of a percent.)
--------------------------------------
Percentage means "per hundred"
---
Set up a proportion:
x/100 = 28/63
x = 100(28/63) = 44.44
=========================
Answer: 44.44%
====================
Cheers,
Stan H.